OpenStudy (anonymous):

please tell me i got this : 10x^6 + 8x^4 ( divided) 2x^2 = 9x ^12

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you familiar with long division?

OpenStudy (anonymous):

o wait. u don't need to. lol

OpenStudy (anonymous):

\[\frac{10x^6 + 8x^4 }{2x^2}=\frac{10x^6}{2x^2}+\frac{8x^4}{2x^2}\] is a start

OpenStudy (anonymous):

divide the numbers, subtract the exponents

OpenStudy (anonymous):

i just don't understnad how u can use equation that fast.

OpenStudy (anonymous):

god-like

OpenStudy (anonymous):

lol practice

OpenStudy (anonymous):

steroids

OpenStudy (anonymous):

those too

OpenStudy (anonymous):

@suzy4321 you got this? divide term by term

OpenStudy (anonymous):

im doing the work il post my answer when im rdy

OpenStudy (anonymous):

ok

OpenStudy (primeralph):

No, no, nooooooooooooooooo 

OpenStudy (primeralph):

You get the point.

OpenStudy (anonymous):

ok i came up with 5x^4 + 4x^2

OpenStudy (primeralph):

Yeah, good job.

OpenStudy (anonymous):

lol nice link primealph

OpenStudy (primeralph):

@chris00 Trust me on that. I go harddddd when I say no.

OpenStudy (anonymous):

yeah, that is it

OpenStudy (anonymous):

good thanks

OpenStudy (anonymous):

LOL